OpenStudy (anonymous):

Solve for the angle θ, where 0≤θ≤2π. cos²(θ)=1/4 θ= ??

myininaya (myininaya):

\[\cos(\theta)=\frac{1}{2} \text{ or } \cos(\theta)=\frac{-1}{2}\]

OpenStudy (hoblos):

+/- π/3

myininaya (myininaya):

your job is to use the unit circle to see when this happens

myininaya (myininaya):

\[\theta=60^o, 300^o , \theta=120^o, 240^o\]

OpenStudy (anonymous):

myininaya where do you see degrees in this problem?

myininaya (myininaya):

it is real easy to convert to degrees if you want it in degrees

OpenStudy (anonymous):

says \[0\leq \theta\leq2\pi\]

myininaya (myininaya):

multiply by pi/180degrees

myininaya (myininaya):

you do know 0 =0 degrees and 2pi=360 degrees

OpenStudy (anonymous):

like hell

myininaya (myininaya):

lol

OpenStudy (anonymous):

\[\sin(x),\cos(x)\] are functions of numbers. put in a number, get out a number.

myininaya (myininaya):

there is no pleasing satellite

OpenStudy (anonymous):

in this case the number is between 0 and 2 pi

myininaya (myininaya):

or 0 and 360 degrees

OpenStudy (anonymous):

hell no

myininaya (myininaya):

hell yes

myininaya (myininaya):

like i said if he wants to convert it to radians he can just multiply by pi/(180degrees)

OpenStudy (anonymous):

if i ask you to solve \[f(x)=x^2+2x=6\] you do not ask me if x is in degrees, radians, Celsius, Fahrenheit, dollars or ruppees x is assumed to be a real number

OpenStudy (anonymous):

use the half angle formula

myininaya (myininaya):

i'm telling the problem what i want to work with when i give the problem my answer i told the problem i want to work in degrees so the problem understood

OpenStudy (anonymous):

sine and cosine are functions of real numbers. as such they correspond to the functions of "angles" only if they are measured in radians

OpenStudy (anonymous):

that does nothing but lead to confusion. convert, solve, convert back. terrible idea \[\frac{\pi}{3}\] is a perfectly good number.

myininaya (myininaya):

so is 60 degrees

OpenStudy (anonymous):

no it is not it is a number

myininaya (myininaya):

what?

OpenStudy (anonymous):



myininaya (myininaya):

i know its a number you aren't making any sense

OpenStudy (anonymous):

sits somewhere to the right of 1 on the number line

myininaya (myininaya):

so are you saying 12 inches and 1 foot don't represent the same measurement?

OpenStudy (anonymous):

it is like saying 0 is 32 because in celsius 0 degrees corresponds to 32 degrees farenheit

OpenStudy (anonymous):

i am saying that sine and cosine are functions of numbers. put in a number, get out a number.

OpenStudy (anonymous):

i agree that 1 foot equals 12 inches, but that doesn't mean that if i have a function \[f(x)=x^2+6x\] then \[f(1)=f(12)\]

myininaya (myininaya):

f(1 foot)=f(12 inches)

OpenStudy (anonymous):

we both agree that the derivative of sine is cosine right?

OpenStudy (anonymous):

functions are functions of real numbers, not if degrees or radians or dollars or feet or yards or parsecs

OpenStudy (anonymous):

what about cos²(θ)=(2+√2) / (4)

OpenStudy (anonymous):

as such, the trig functions correspond the the high school function of angles only if the angles are measured in radians

OpenStudy (anonymous):

this one is a pain gimmick is to rewrite \[\cos^2(x)\] as \[\frac{1}{2}(\cos(2x)+1)\]

OpenStudy (anonymous):

that will give you \[\frac{1}{2}(\cos(2x)+1)=\frac{2+\sqrt{2}}{4}\] \[2(\cos(2x)+1)=2+\sqrt{2}\] \[2\cos(2x)=\sqrt{2}\] \[\cos(2x)=\frac{\sqrt{2}}{2}\] and so \[2x=\frac{\pi}{4}\] or \[2x=\frac{7\pi}{4}\] etc

OpenStudy (anonymous):

shall i divide 2x by 2??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so it will be 7π/8 ,π4/8 and what are the other 2 angles